# Generative Motion Stylization within Canonical Motion Space

### v1

@@ Abstract.
Stylized motion breathes life into characters. However, the fixed skeleton structure and style representation hinder existing data-driven motion synthesis methods from generating stylized motion for various characters. In this work, we propose a generative motion stylization pipeline, named Motion, for synthesizing diverse and stylized motion on cross-structure characters using cross-modality style prompts. Our key insight is to embed motion style into a cross-modality latent space and perceive the cross-structure skeleton topologies, allowing for motion stylization within a canonical motion space. Specifically, the large-scale Contrastive-Language-Image-Pre-training (CLIP) model is leveraged to construct the cross-modality latent space, enabling flexible style representation within this space. Additionally, two topology-encoded tokens are learned to capture the canonical and specific skeleton topologies, facilitating cross-structure topology shifting. Subsequently, the topology-shifted stylization diffusion is designed to generate motion content for the specific skeleton and stylize it in the shifted canonical motion space using multi-modality style descriptions. Through an extensive set of examples, we demonstrate the flexibility and generalizability of our pipeline across various characters and style descriptions. Qualitative and quantitative experiments underscore the superiority of our pipeline over state-of-the-art methods, consistently delivering high-quality stylized motion across a broad spectrum of skeletal structures.
@@ 1. Introduction
Style serves as a pivotal element in animations, mirroring diverse facets of a character, including personality, habits, emotions, and intentions. Therefore, the expressiveness of style plays a crucial role in motion synthesis and enriches storytelling in computer animation. However, the complexity and time-consuming nature of the traditional motion stylization process presents significant barriers to entry (Jang et al., 2022). In recent years, learning-based motion synthesis methods have emerged in the community (Wan et al., 2023; Dou et al., 2023). Still, they are often constrained to a fixed skeleton structure, generating motion content without expressing the nuances of motion styles. Thus, they fall short of synthesizing stylized motion for a wide variety of characters.
@@
By delving deeper into the underlying reasons, we have identified two primary challenges impeding the development of a comprehensive stylized motion synthesis pipeline. The first challenge involves expressing motion style on cross-structure skeletons. Existing methods (Aberman et al., 2020b; Dong et al., 2020; Park et al., 2021) for motion style transfer often oversimplify this obstacle by assuming a fixed skeletal structure between the source and target characters, thereby neglecting the intricacies of the characters’ skeletal topologies. This inherent assumption significantly limits their applicability for transferring motion style among characters with disparate skeletal structures. Moreover, the predominant focus of existing stylized motion datasets (Xia et al., 2015; Mason et al., 2022) is on standard human skeletons, making it difficult for learning-based methods to stylize non-standard skeletal motions. Consequently, perceiving the skeleton topology and effectively expressing motion style on cross-structure skeletons emerge as a critical challenge.
@@
The second challenge pertains to the utilization of cross-modality style representations. Existing motion generation methods (Guo et al., 2020; Jiang et al., 2023) have explored various conditions to control the content of the generated motion. However, these methods struggle to effectively control the motion style through these multi-modality conditions, resulting in generated motion that is often perceived as dull and monotonous. Additionally, previous motion style transfer methods (Smith et al., 2019; Song et al., 2023) are constrained to using a motion sequence or a category label as the style description, rendering flexible style control impossible. Thus, the incorporation of cross-modality style representations for more flexible and user-friendly style editing stands out as another significant challenge.
@@
To address these challenges, we introduce a new generative motion stylization pipeline, Motion, capable of synthesizing diverse motion for a wide range of characters and stylizing it using motion sequences, text, images, or videos as style descriptions, as Figure 1 shows. In Motion, two novel techniques, i.e. cross-modality style embedding and cross-structure topology shifting, are proposed to construct a canonical motion space. In this space, motion content from various skeletons can be stylized using the extracted style embedding by adjusting the mean and variance of the generated motion features. With these two key techniques, the topology-shifted stylization diffusion is presented to synthesize diverse and stylized motion, establishing a more accessible way of animation creation.
@@
Cross-modality style embedding is to embed style descriptions in various formats into the canonical motion space. Drawing inspiration from GestureDiffuCLIP (Ao et al., 2023), which employs CLIP (Radford et al., 2021) latents for the automatic creation of stylized gestures, we align the motion features of the SMPL skeleton (Loper et al., 2015) with CLIP latents and construct a shared space that represents motion style as adaptive mean and variance applied to the motion features. The multi-modality style embedding in the shared latent space enables flexible style descriptions for motion stylization in our Motion.
@@
Cross-structure topology shifting is to transfer motion features between canonical and specific motion spaces. To achieve this, we utilize two learnable topology-encoded tokens (TET) for capturing the canonical and specific skeleton topologies. Each TET is followed by a graph convolutional layer (GCL) with a pre-defined adjacency matrix, serving as a topology prior of the skeletal structure. Subsequently, the motion space is shifted using the cross-attention mechanism. This topology shifting strategy enables cross-structure motion stylization in our Motion.
@@
In response to the fact posed by the scarcity motion sequences of arbitrary skeletal structure, we designed our Motion based on the SinMDM (Raab et al., 2023). SinMDM requires only one motion sequence to learn the motion content and can synthesize variable-length motion sequences that retain the core motion elements of the single input. It is worth noting that, distinct from SinMDM, the focus of this work is on cross-structure motion stylization and the utilization of cross-modality style representations.
@@
We introduce Motion, a generative motion stylization pipeline that, to our knowledge, marks the first attempt to synthesize cross-structure diverse and stylized motion using cross-modality style descriptions.
We propose two novel techniques in Motion i.e., cross-modality style embedding and cross-structure topology shifting. These strategies collaborate to construct a canonical motion space, enabling skeletal topology perception and flexible style representations for motion stylization.
Extensive experiments demonstrate the effectiveness and generalizability of our Motion qualitatively and quantitatively, achieving higher motion quality and style diversity in comparison with existing learning-based methods.
@@ 2. Related Work
Motion stylization has been a longstanding challenge within the realm of computer animation (Amaya et al., 1996; Brand and Hertzmann, 2000; Hsu et al., 2005). Recently, learning-based approaches (Holden et al., 2016, 2017; Du et al., 2019; Park et al., 2021; Jang et al., 2023) started sparkling in the community. For example, Xia et al. (2015) proposed an online learning algorithm to capture complex relationships between pairwise motion styles. Yumer and Mitra (2016) represented motion style in the frequency domain, extracting style features without the need for spatial matching. Smith et al. (2019) presented a neural network-based style transfer model, enabling the adjustment of output motion in a latent space. These methods highlight the research trend of achieving flexibility and diversity in style control. However, they are limited to representing labeled or sample-based motion styles, lacking support for multi-modality style descriptions.
@@
Motion generation has emerged as a central focus of animation creation in recent years, propelled by advancements in generative deep models (Goodfellow et al., 2014; Ho et al., 2020). Recent cutting-edge methods often explore various conditions for human motion generation, which include but are not limited to action labels (Guo et al., 2020; Petrovich et al., 2021), audio (Ao et al., 2022), and text (Wan et al., 2023). For example, Tevet et al. (2022b) introduced the Motion Diffusion Model (MDM), aiming for natural and expressive human motion generation from text prompts. Building upon MDM, Chen et al. (2023) employed a Variational Auto-encoder (VAE) to enhance motion representation, and Zhang et al. (2023a) integrated a retrieval mechanism to refine the denoising process. Additionally, Zhang et al. (2023b) proposed a text-driven animation pipeline (TapMo) for generating motion in a broad spectrum of skeleton-free characters. However, these methods typically require large amounts of data for training and overlook motion style in the generation process. In our work, recognizing the scarcity of motion data for non-standard skeletons, we draw inspiration from SinMDM (Raab et al., 2023) to construct a generative model learning from a single motion sequence. More importantly, our pipeline can control the style of cross-structure skeletal motion through multi-modality style descriptions in a learned canonical motion space.
@@ 3. Method
Subsequently, the diffusion model takes the noised motion and the step as inputs to generate the stylized motion through denoising steps. All motions in this work are represented using the 6D rotation features proposed by Zhou et al. (2019). Throughout this process, the style feature is equipped to provide fine-grained dynamic information about the motion style. The style embedding influences the motion feature in canonical space using the AdaIN strategy (Huang and Belongie, 2017), controlling the style property. Notably, two TETs, i.e., and , are employed to shift the motion feature space between canonical and specific settings. This is achieved through the cross-attention mechanism used in the style decoder and the diffusion decoder, allowing the model to perceive the skeletal topologies and control motion style on cross-structure skeletons. This inverse diffuse process is formulated as:
@@ 3.1. Cross-modality Style Embedding
As depicted in Figure 3, drawing inspiration from MotionCLIP (Tevet et al., 2022a), we design a transformer-based motion encoder to embed into a style feature and a content feature , where is the time dimension and is the channel. The style embedding and are then obtained by mapping through a linear layer. Next, we repeat along the temporal dimension and concatenate it with to construct the input motion feature of dimensions for the style decoder. Subsequently, the transformer-based style decoder utilizes the input motion feature as the Key and Value of the cross-attention mechanism. Meanwhile, it takes the GCL refined canonical TET as the Query to map the motion feature from to dimensions. At this point, the motion feature and the style feature are all expressed in a canonical motion space. This process is formulated as:
where and denote the style decoder and its learnable parameter, respectively. Finally, the AdaIN and a linear layer are applied to reconstruct the input motion using , , and from the canonical motion space.
@@
With these two losses, the style encoder and the style decoder can be trained by:
@@ 3.2. Cross-structure Topology Shifting
To enable general motion stylization beyond the standard human skeleton, we introduce the cross-structure topology shifting strategy, which involves transferring the skeletal topology of the motion in the latent feature space. As illustrated in the right part of Figure 4, for a type of skeleton with joints, we exploit a learnable topology-encoded token to capture the structure and order of the skeleton joints. Leveraging the graph convolutional layer, which incorporates a strict topology prior, we input the initial TET into a GCL to refine the encoded topology, thus facilitating the learning of the skeletal structure. This process can be formulated as:
@@ 3.3. Topology-shifted Stylization Diffusion
Specifically, TSD comprises three network modules: the diffusion encoder, the style decoder, and the diffusion decoder. First, the QnA-based diffusion encoder takes the noised motion and the step as inputs, producing the motion content feature . Simultaneously, the style encoder extracts the style feature from an SMPL motion sequence. Then, following a process similar to the cross-modality style embedding shown in Figure 3, the style feature is repeated, concatenated with the motion feature, and fed into the pre-trained style decoder. At this stage, the style decoder, integrated with the canonical TET, transfers the style feature to the canonical motion space, allowing for stylization using the style embedding. Notably, the parameters of the style encoder, style decoder, and canonical TET are all pre-trained and frozen in this process.
@@
The style loss aims to guarantee that the stylized canonical motion feature embodies the corresponding motion style present in the style prompt motion . To achieve this, we employ the final linear layer of the pre-trained style decoder to reconstruct into a motion sequence compatible with the standard SMPL skeleton. Subsequently, we extract its style feature using the pre-trained style encoder. The style loss is then calculated as:
@@ 4. Experiments
Implementation details. We implement our pipeline using PyTorch framework(Paszke et al., 2019). In the cross-modality style embedding (refer to Section 3.1), the motion style encoder comprises a four-layer transformer-encoder, while the style decoder is constructed with a four-layer transformer-decoder. In TSD (refer to Section 3.3), the diffusion encoder consists of an eight-layer QnA-encoder, and the structure of the diffusion decoder mirrors that of the style decoder. In our pipeline, the feature channel size is set to . The number of the canonical joints is , representing the SMPL skeleton joints excluding the root joint. For video prompts, we randomly sample five frames to extract style features, which are computed as the average image features of the sampled frames. Please see the supplementary material for more details.
@@
| Content | Style | Glo-D | Loc-D | Content | Style | Glo-D | Loc-D | |
@@ 4.1. Qualitative Results
Comparison with the baselines. Figure 5 illustrates the qualitative comparison between our Motion and SinMDM. In their study, the authors trained the model on style motion and fused content motion during the inference process for stylization. However, we observed that this approach often results in generated outputs closely resembling the style motion, potentially neglecting the preservation of content motion, as shown in the fourth column of Figure 5. Thus, we additionally implemented SinMDM by training the model on the source motion instead of the style motion.
Our Motion effectively handles various skeleton structures, stylizing the source motion based on key features of multi-modality style prompts. Importantly, our Motion well preserves the motion content in the generated results. For instance, when using a walking motion with arms behind the back as the style prompt and a jumping motion as the source, our result maintains the jumping content and adjusts the skeletal arms to resemble the stylized motion. In contrast, SinMDM tends to lose the jumping content, mistakenly generating a walking motion. Furthermore, SinMDM fails to transfer the style to the content motion. In addition, when the skeletal structure is complex and there is a significant difference between the style prompt and source motion, both SinMDM and SinMDM face mode collapse, resulting in static motion sequences, as depicted in the third row of Figure 5. In summary, Motion inherits the strengths of SinMDM in generating diverse sequences from a single source. Additionally, Motion is capable of learning skeletal structure and utilizing multi-modality cues for motion stylization.
@@
The concept of cross-modality style embedding is inspired by MotionCLIP, with a key distinction being that our approach independently learns the motion content feature and the motion style feature, reconstructing the input motion using both features. To assess this design, we conducted an experiment by employing MotionCLIP as the style encoder and incorporating a learnable style decoder in TSD. The results, as shown in Figure 7, indicate that Motion (CLIP) fails to reconstruct the dynamic source content, yielding a static stylized result. Additionally, in Motion, we concatenate the temporally repeated style feature with the content feature to obtain the motion feature, as depicted in Figure 3 and Figure 4. This design is crucial for learning dynamic motion styles. The result of Motion (w/o ) demonstrates that without this process, the model fails to express the motion style in the generated motion.
The reason for the aforementioned failed result is that the motion style features aligned to the CLIP space only contain static posture information, posing a challenge for the model to transfer dynamic motion styles to the source motion. Thanks to our approach, which involves decoupled learning of motion content and style features, along with the weights transfer strategy of the pre-trained style decoder, TSD in our Motion can be theoretically viewed as comprising two complementary diffusion processes. Initially, it leverages the style feature as a condition to generate the dynamic content of the style motion and subsequently stylizes the source motion in the canonical motion space. Consequently, our cross-modality style embedding plays a pivotal role in achieving flexible motion stylization in the generative process.

### v2

# Generative Motion Stylization of Cross-structure Characters within Canonical Motion Space

@@ Abstract.
Stylized motion breathes life into characters. However, the fixed skeleton structure and style representation hinder existing data-driven motion synthesis methods from generating stylized motion for various characters. In this work, we propose a generative motion stylization pipeline, named Motion, for synthesizing diverse and stylized motion on cross-structure characters using cross-modality style prompts. Our key insight is to embed motion style into a cross-modality latent space and perceive the cross-structure skeleton topologies, allowing for motion stylization within a canonical motion space. Specifically, the large-scale Contrastive-Language-Image-Pre-training (CLIP) model is leveraged to construct the cross-modality latent space, enabling flexible style representation within it. Additionally, two topology-encoded tokens are learned to capture the canonical and specific skeleton topologies, facilitating cross-structure topology shifting. Subsequently, the topology-shifted stylization diffusion is designed to generate motion content for the particular skeleton and stylize it in the shifted canonical motion space using multi-modality style descriptions. Through an extensive set of examples, we demonstrate the flexibility and generalizability of our pipeline across various characters and style descriptions. Qualitative and quantitative comparisons show the superiority of our pipeline over state-of-the-arts, consistently delivering high-quality stylized motion across a broad spectrum of skeletal structures.
@@ 1. Introduction
Style serves as a pivotal element in animations, mirroring diverse facets of a character, including personality, habits, emotions, and intentions. Therefore, the expressiveness of motion style plays a crucial role in motion synthesis and enriches storytelling in computer animation. However, the complexity and time-consuming nature of the traditional motion stylization process presents significant barriers to entry (Jang et al., 2022). In recent years, learning-based motion synthesis methods have emerged in the community (Wan et al., 2023; Dou et al., 2023). Still, they are often constrained to a fixed skeleton structure, generating motion content without expressing the nuances of motion styles. Thus, they fall short of synthesizing stylized motion for a wide variety of characters.
@@
By delving deeper into the underlying reasons, we have identified two primary challenges impeding the development of a comprehensive stylized motion synthesis. The first challenge involves expressing motion style on cross-structure skeletons. Existing methods (Aberman et al., 2020b; Dong et al., 2020; Park et al., 2021) for motion style transfer oversimplify this obstacle by assuming a fixed skeletal structure between the source and target characters, thereby neglecting the intricacies of the characters’ skeletal topologies. This inherent assumption significantly limits their applicability for transferring motion style among characters with disparate skeletal structures. Moreover, the predominant focus of existing stylized motion datasets (Xia et al., 2015; Mason et al., 2022) is on standard human skeletons, making it difficult for learning-based methods to stylize non-standard skeletal motions. Consequently, perceiving the skeleton topology and effectively expressing motion style on cross-structure skeletons emerge as a critical task.
@@
The second challenge pertains to the utilization of cross-modality style representations. Existing motion generation methods (Guo et al., 2020; Jiang et al., 2024) have explored various conditions to control the content of the generated motion. However, these methods struggle to effectively control the motion style through these multi-modality conditions, resulting in the generated motion often perceived as dull and monotonous. Additionally, the previous motion style transfer methods (Smith et al., 2019; Song et al., 2023) are constrained to using motion sequences or category labels as the style description, rendering flexible style control impossible. Thus, incorporating cross-modality style representations for more flexible and user-friendly motion-style editing stands out as another significant task.
@@
To address these challenges, we introduce a new generative motion stylization pipeline, Motion, capable of synthesizing diverse motion for a wide range of characters and stylizing it using motion sequences, text, images, or videos as style descriptions, as Figure 1 shows. In Motion, two novel techniques, i.e. cross-modality style embedding and cross-structure topology shifting, are explored to construct a canonical motion space. In this space, motion content from various skeletons can be stylized using the extracted style embedding by adjusting the mean and variance of the generated motion features. With these two key techniques, the topology-shifted stylization diffusion is presented to synthesize diverse and stylized motion, establishing a more accessible way of animation creation.
@@
Cross-modality style embedding is to embed style descriptions in various formats into the canonical motion space. Drawing inspiration from GestureDiffuCLIP (Ao et al., 2023), which employs CLIP latents (Radford et al., 2021) for the automatic creation of stylized gestures, we align the motion features of the SMPL skeleton (Loper et al., 2015) with CLIP latents and construct a shared space that represents motion style as adaptive mean and variance applied to the motion features. The multi-modality style embedding in the shared latent space enables flexible style descriptions for motion stylization in our Motion.
@@
Cross-structure topology shifting aims to transfer motion features between canonical and specific motion spaces. To achieve this, we utilize two learnable topology-encoded tokens (TET) for capturing the canonical and specific skeleton topologies. Each TET is followed by a graph convolutional layer (GCL) with a pre-defined adjacency matrix, serving as a topology prior of the skeletal structure. Subsequently, the motion space is shifted using the cross-attention mechanism. This topology shifting strategy enables cross-structure motion stylization in our Motion.
@@
In response to the fact posed by the scarcity motion sequences of arbitrary skeletal structure, we designed our Motion based on the SinMDM (Raab et al., 2023). SinMDM requires only one motion sequence to learn the motion content and can synthesize variable-length motion sequences that retain the core motion elements of the single input. Remarkably, distinct from SinMDM, the focus of this work is on the stylization of cross-structure motion and the utilization of cross-modality style representations.
@@
We propose Motion, a generative motion stylization pipeline that, to our knowledge, marks the pioneering attempt to synthesize cross-structure diverse and stylized motion by utilizing cross-modality style descriptions.
We exploit two novel techniques in Motion i.e., cross-modality style embedding and cross-structure topology shifting. These techniques collaborate to construct a canonical motion space, enabling skeletal topology perception and flexible style representations for motion stylization.
Extensive experiments demonstrate the effectiveness and generalizability of our Motion, which achieves higher motion quality and style diversity in comparison with the prior learning-based methods.
@@ 2. Related Work
Motion stylization has been a longstanding challenge within the realm of computer animation (Amaya et al., 1996; Brand and Hertzmann, 2000; Hsu et al., 2005; Zhang et al., 2023b). Recently, deep learning-based approaches (Holden et al., 2016, 2017; Du et al., 2019; Park et al., 2021; Jang et al., 2023) started sparkling in the community. For example, Xia et al. (2015) proposed an online learning algorithm to capture complex relationships between pairwise motion styles. Yumer and Mitra (2016) represented motion style in the frequency domain, extracting style features without the need for spatial matching. Smith et al. (2019) presented a neural network-based style transfer model, enabling the adjustment of output motion in a latent space. These methods highlight the research trend of achieving flexibility and diversity in style control. However, they are limited to representing labeled or sample-based motion styles, lacking support for multi-modality style descriptions.
@@
Motion generation has emerged as a central focus of animation creation in recent years, propelled by advancements in generative deep models (Goodfellow et al., 2014; Ho et al., 2020). Recent cutting-edge methods often explore various conditions for human motion generation, which include but are not limited to action labels (Guo et al., 2020; Petrovich et al., 2021), audio (Ao et al., 2022), and text (Wan et al., 2023). For example, Tevet et al. (2023) introduced the Motion Diffusion Model (MDM), aiming for natural and expressive human motion generation from text prompts. Building upon MDM, Chen et al. (2023) employed a Variational Auto-encoder (VAE) to enhance motion representation, and Zhang et al. (2023a) integrated a retrieval mechanism to refine the denoising process. Additionally, Zhang et al. (2024) proposed a text-driven animation pipeline (TapMo) for generating motion in a broad spectrum of skeleton-free characters. However, these methods typically require large amounts of data for training and overlook motion style in the generation process. In our work, recognizing the scarcity of motion data for non-standard skeletons, we draw inspiration from SinMDM (Raab et al., 2023) to construct a generative model learning from a single motion sequence. More importantly, our pipeline can control the style of cross-structure skeletal motion through multi-modality style descriptions in a learned canonical motion space.
@@ 3. Method
Subsequently, the diffusion model takes the noised motion and the step as inputs to generate the stylized motion through denoising steps. All motions in this work are represented using the 6D rotation features proposed by Zhou et al. (2019). Throughout this process, the style feature is equipped to provide fine-grained dynamic information about the motion style. The style embedding influences the motion feature in canonical space using the AdaIN strategy (Huang and Belongie, 2017), controlling the style property. Notably, two TETs, i.e., and , are employed to shift the motion feature space between canonical and specific settings. This is achieved through the cross-attention mechanism used in the canonical decoder and the diffusion decoder, allowing the model to perceive the skeletal topologies and control motion style on cross-structure skeletons. This inverse diffuse process is formulated as:
@@ 3.1. Cross-modality Style Embedding
As depicted in Figure 3, drawing inspiration from MotionCLIP (Tevet et al., 2022), we design a transformer-based motion encoder to embed into a style feature and a content feature , where is the time dimension and is the channel. The style embedding and are then obtained by mapping through a linear layer. Next, we repeat along the temporal dimension and concatenate it with to construct the input motion feature of dimensions for the canonical decoder. Subsequently, the transformer-based canonical decoder utilizes the input motion feature as the Key and Value of the cross-attention mechanism. Meanwhile, it takes the GCL refined canonical TET as the Query to map the motion feature from to dimensions. At this point, the motion feature and the style feature are all expressed in a canonical motion space. This process is formulated as:
where and denote the canonical decoder and its learnable parameter, respectively. Finally, the AdaIN and a linear layer are applied to reconstruct the input motion using , , and from the canonical motion space.
@@
With these two losses, the style encoder and the canonical decoder can be trained by:
@@ 3.2. Cross-structure Topology Shifting
To enable general motion stylization beyond the standard human skeleton, we introduce the cross-structure topology shifting strategy, which involves transferring the skeletal topology of the motion in the latent feature space. As illustrated in the right part of Figure 4, for a type of skeleton B with joints, we exploit a learnable topology-encoded token to capture the structure and order of the skeleton joints. Leveraging the graph convolutional layer inspired by SAME (Lee et al., 2023), which incorporates a strict topology prior, we input the initial TET into a GCL to refine the encoded topology, thus facilitating the learning of the skeletal structure. This process can be formulated as:
@@ 3.3. Topology-shifted Stylization Diffusion
Specifically, TSD comprises three network modules: the diffusion encoder, the canonical decoder, and the diffusion decoder. First, the QnA-based diffusion encoder takes the noised motion and the step as inputs, producing the motion content feature . Simultaneously, the style encoder extracts the style feature from an SMPL motion sequence. Then, following a process similar to the cross-modality style embedding shown in Figure 3, the style feature is repeated, concatenated with the motion feature, and fed into the pre-trained canonical decoder. At this stage, the canonical decoder, integrated with the canonical TET, transfers the style feature to the canonical motion space, allowing for stylization using the style embedding. Notably, the parameters of the style encoder, canonical decoder, and canonical TET are all frozen in this process.
@@
The style loss aims to guarantee that the stylized canonical motion feature embodies the corresponding motion style present in the style prompt motion . To achieve this, we employ the final linear layer of the pre-trained canonical decoder to reconstruct into a motion sequence compatible with the standard SMPL skeleton. Subsequently, we extract its style feature using the pre-trained style encoder. The style loss is then calculated as:
@@ 4. Experiments
| Content↓ | Style↓ | Glo-D→ | Loc-D→ | Content↓ | Style↓ | Glo-D→ | Loc-D→ | |
@@
Implementation details. We implement our pipeline using PyTorch framework(Paszke et al., 2019). In the cross-modality style embedding (refer to Section 3.1), the motion style encoder comprises a four-layer transformer-encoder, while the canonical decoder is constructed with a four-layer transformer-decoder. In TSD (refer to Section 3.3), the diffusion encoder consists of an eight-layer QnA-encoder, and the structure of the diffusion decoder mirrors that of the canonical decoder. In our pipeline, the feature channel size is set to . The number of the canonical joints is , representing the SMPL skeleton joints excluding the root joint. For video prompts, we randomly sample five frames to extract style features, which are computed as the average image features of the sampled frames. Please see the supplementary material for more details.
@@ 4.1. Qualitative Results
Comparison with the baselines. Figure 5 illustrates the qualitative comparison between our Motion and SinMDM. In their study, the authors trained the model on style motion and fused content motion during the inference process for stylization. However, we observed that this approach often results in generated outputs closely resembling the style motion, potentially neglecting the preservation of content motion, as shown in the fourth column of Figure 5. Thus, we additionally implemented SinMDM∗ by training the model on the source motion instead of the style motion.
Our Motion effectively handles various skeleton structures, stylizing the source motion based on key features of multi-modality style prompts. Importantly, our Motion well preserves the motion content in the generated results. For instance, when using a walking motion with arms behind the back as the style prompt and a jumping motion as the source, our result maintains the jumping content and adjusts the skeletal arms to resemble the stylized motion. In contrast, SinMDM tends to lose the jumping content, mistakenly generating a walking motion. Furthermore, SinMDM∗ fails to transfer the style to the content motion. In addition, when the skeletal structure is complex and there is a significant difference between the style prompt and source motion, both SinMDM and SinMDM∗ face mode collapse, resulting in static motion sequences, as depicted in the third row of Figure 5. In summary, Motion inherits the strengths of SinMDM in generating diverse sequences from a single source. Additionally, Motion is capable of learning skeletal structure and utilizing multi-modality cues for motion stylization.
@@
The concept of cross-modality style embedding is inspired by MotionCLIP, with a key distinction being that our approach independently learns the motion content feature and the motion style feature, reconstructing the input motion using both features. To assess this design, we conducted an experiment by employing MotionCLIP as the style encoder and incorporating a learnable canonical decoder in TSD. The results, as shown in Figure 7, indicate that Motion (CLIP) fails to reconstruct the dynamic source content, yielding a static stylized result. Additionally, in Motion, we concatenate the temporally repeated style feature with the content feature to obtain the motion feature, as depicted in Figure 3 and Figure 4. This design is crucial for learning dynamic motion styles. The result of Motion (w/o ) demonstrates that without this process, the model fails to express the motion style in the generated motion.
The reason for the aforementioned failed result is that the motion style features aligned to the CLIP space only contain static posture information, posing a challenge for the model to transfer dynamic motion styles to the source motion. Thanks to our approach, which involves decoupled learning of motion content and style features, along with the weights transfer strategy of the pre-trained canonical decoder, TSD in our Motion can be theoretically viewed as comprising two complementary diffusion processes. Initially, it leverages the style feature as a condition to generate the dynamic content of the style motion and subsequently stylizes the source motion in the canonical motion space. Consequently, our cross-modality style embedding plays a pivotal role in achieving flexible motion stylization in the generative process.